OpenStudy (anonymous):

find the equation of the line having the given slope and containing the given point m=7/8(7,-5)

OpenStudy (anonymous):

with the slope 7/8, we have the equation, y = 7/8 x + c or, 8y = 7x + 8c ............................ (1) Now, equation (1) satisfies the point (7, -5) So, we get, -40 = 49 + 8c or, c = - 89/8 Putting the value of c in (1), 8y = 7x - 8*89/8 or, 8y = 7x - 89 or, 7x - 8y - 89 = 0 Hence is the equation.